# A Controllable and Realistic Framework for Evaluating Microservice Scheduling in Cloud-Edge Continuum

### v2

@@ 5.1.1. Istio Service Mesh Implementation
To achieve fine-grained observation into bidirectional traffic between UM and DM microservices, including their replicas, we implemented the UM-DM Traffic Profiler using the Istio service mesh [^1] Istio is an open-source service mesh employing sidecar containers to facilitate secure communication, traffic management, and load balancing. (Istio: An open source service mesh, 2024). The service mesh operates as an additional infrastructure layer integrated into containerized microservice deployments. With Istio deployed, all microservice traffic is transparently proxied via injected sidecar containers.
@@
In Kubernetes clusters with Istio service mesh enabled, each microservice pod includes the application container alongside an Envoy sidecar container. As illustrated in Figure 5, the service mesh architecture proxies traffic among microservice containers. In the Data Plane, orange rectangles represent the microservice containers, while green rectangles denote the sidecar containers that manage traffic and maintain dependencies between services. The green lines indicate data flows and dependency relationships among microservices. In the Control Plane, the Istio service mesh manages services and interactions via components including Pilot, Mixer and Citadel for different purposes.
@@
Figure 5. Overview of a service mesh structure consisting of data plane and control plane.
5.1.2. Overhead Analysis
@@
Deploying a service mesh introduces additional processing delays due to the sidecar proxy layer. However, empirical evaluations indicate these delays minimally impact overall latency performance. According to (Istio: Performance and Scalability, 2024), in Istio 1.21.2 with telemetry v2 enabled, each request passing through client-side and server-side Envoy proxies incurs a latency increase of approximately 0.18 milliseconds at the 90th percentile and 0.25 milliseconds at the 99th percentile compared to baseline performance without proxies. These results, obtained using a 1kB payload at 1000 requests per second across varying client connections (2, 4, 8, 16, 32, 64), were measured at the CNCF Community Infrastructure Lab (CNCF Community Infrastructure Lab : An on-demand infrastructure resource for CNCF developers.(2024), CIL). Thus, the added latency overhead from the service mesh is negligible and acceptable for most practical scenarios.
5.2. Call-graph Builder
We introduce the following terminology:
(1) $\text{stress}^{\mu}_{\sigma}(\mu^{UM},\sigma^{DM},\Delta t)=\frac{\text{Bi-directional\_traffic}(\mu^{UM},\sigma^{DM})}{2\Delta t}$
[…]
Algorithm 1 constructs the call-graphs with dependency and associated traffic load based on the defined stress elements. Initially, the algorithm retrieves all running microservices deployed within a namespace, which is used to seperate different applications. It then initializes a directed graph, with each microservice represented as a graph node. For each microservice pair, the algorithm calculates the stress using real-time traffic metrics collected over the specified interval. If the calculated traffic stress is greater than zero, an edge representing the dependency and corresponding traffic load is created. The constructed graph thus reflects both the dependency topology and associated traffic stress, enabling accurate evaluation of scheduling decisions.
@@ 5.2. Call-graph Builder
2:Output: A directed call-graph $G$ with traffic edges
@@
Figure 6. (a) The process of enqueueing packets and dequeuing to different destinations depending on classful disciplines. (b) Packets classified for destination-specific delay emulation. (c) Packets classified for bandwidth shaping towards different destinations.
6. Design of Networking Dynamics Manager
The Networking Dynamics Manager in the proposed framework iDynamics consists of two main components: the Emulator and Measurer. In cloud-edge computing, accurately emulating dynamic network conditions such as latency variations and bandwidth fluctuations is critical for evaluating scheduling policies against SLA targets. However, performing these evaluations directly in production environments is challenging and inefficient. iDynamics addresses this issue by providing a controllable networking dynamics emulator that accurately replicates delays and bandwidths. Moreover, precise and efficient measurement of these emulated dynamics is essential for reliable evaluation.
6.1. Linux Traffic Control
@@
u32 filter: Filters in tc play a crucial role in directing traffic to specific classes or queues by identifying and matching packet characteristics. u32 filter is one of the most versatile and widely used filters in tc. It allows users to match arbitrary bits of a packet header, enabling highly specific rules. For example, users can match fields like source/destination IPs, port numbers, and protocol types by extracting specific bits from the header.
For the dynamic network emulations in iDynamics, classful qdiscs and u32 filter are implemented to meet the goals of changing cross-node communication delays and bandwidths in the cloud-edge continuum. Moreover, after injecting the dynamic delays or shaping available bandwidths, Cross-node Measurer is designed to measure these delays and bandwidths in the current cloud-edge continuum because the accurate and efficient measurement of these dynamics is crucial for designing and evaluating networking-related scheduling policies.
6.2. Emulator: A classful qdisc design
To address these limitations, we propose a customized delay injection scheme that classifies packets using filters to distinguish egress packets and direct them based on their IP destinations (see Figure 6(b)). Additionally, we reserve an extra channel for default packet transmission without injected delays, ensuring that the performance of other services remains unaffected, for example, the packets from/to the master node and outside internet services like Google. We implemented this scheme using the tc networking tool and the Hierarchical Token Bucket (htb).
Furthermore, we employ the technique illustrated in Figure 6 to design an algorithm for emulating customized cross-node delays, as presented in Algorithm 2. To simulate cross-node delays more realistically, we incorporate several factors: the base latency bl (representing the ideal minimal latency); the maximum additional latency mal; the distance factor df (accounting for latency due to physical separation); and the congestion factor cf (which simulates network uncertainties induced by traffic congestion). Here, $\textsc{RandUni}(0,mal)$ denotes the additional delay generated according to a random uniform distribution between 00 and $mal$. Thus, the emulated communication delay from node $i$ to node $j$ in a cluster of $N$ cloud-edge nodes is given by the following equation:
(2) $\text{delay}^{j}_{i}=\left\lfloor\left(bl+\textsc{RandUni}(0,mal)\times\frac{\lvert i-j\rvert}{N}\right)\times cf\right\rfloor$
@@ 6.2. Emulator: A classful qdisc design
Emulation for Customized Bandwidths: Shaping customized available bandwidths between different cloud-edge node pairs is also crucial for creating dynamic networking conditions, thereby enhancing the generalizability of evaluated scheduling policies. Prior research—such as the work presented in (Gouveia et al., 2020) and (Liechti et al., 2019)—has demonstrated the feasibility of emulating dynamic network conditions, including bandwidth variations, using decentralized emulation techniques. However, these approaches typically apply a global or node-level configuration that does not allow fine-grained, per-destination control.
In contrast, our proposed emulation scheme leverages the classful qdisc with unique traffic classes and utilizes the u32 filter to match different IP destinations. As illustrated in Figure 6(c), this approach enables the emulation of distinct bandwidth settings for different destination nodes, even when originating from the same source cluster node. This fine-grained control is essential for accurately emulating heterogeneous network conditions in cloud-edge environments.
6.3. Measurer: Distributed Agent Design
In practical cloud-edge computing environments, fluctuating network conditions—including varying delays and bandwidths between nodes—significantly impact microservice performance and SLA compliance. High latency between cluster nodes negatively affects microservice communication, while dynamically changing bandwidth can lead to congestion, increased packet loss, and potential SLA violations. To accurately capture these dynamic cross-node conditions, we introduce the Measurer, a distributed agent-based measurement module within the Networking Dynamics Manager component of iDynamics (Figure 4).
@@ 6.3. Measurer: Distributed Agent Design
Figure 7. Design of Measurer for delay and bandwidth measurement through distributed agents across cloud-edge nodes
@@
Design Overview: The Measurer uses a unified approach for capturing cross-node delays and bandwidths through distributed measuring agents. Although both delay and bandwidth measurements follow a similar structure, we illustrate the design using delay measurement for clarity. The module consists of a centralized information processing unit and distributed lightweight agents. The centralized processing unit maintains minimal connections with agents, aggregating the collected data efficiently.
@@
The distributed agents run as lightweight containers deployed across cluster nodes, dedicated to measuring and reporting network metrics. To ensure robustness and adaptability during cluster scaling, we implement an automatic scaling mechanism for these agents. When new nodes join the cluster, corresponding agents are automatically instantiated, ensuring consistent and accurate measurements. Conversely, when nodes are removed, their associated agents are gracefully terminated.
@@
Overhead Analysis: Regarding the overhead of measurement in a real cloud-edge cluster, we optimized the measurement using parallel processing. The measurement tasks for delays and bandwidths are designed to run as concurrent tasks, and each of the measured results is aggregated into a result dictionary respectively. Additionally, the distributed delay measuring agents are lightweight (about 0.2 MiB) and stable (running over 6 months without any failure). Each node runs a delay-measuring agent developed from the curlimages/curl image, consuming around 0.2 MiB of memory per node. Similarly, each node also runs a bandwidth-measuring agent implemented from the networkstatic/iperf3 image, consuming around 0.84 MiB of memory on each node. As the measurement tasks are designed to run in parallel as well, the measuring speed can be adaptively tuned by increasing or decreasing the concurrency, depending on the current load levels in the cluster.
7. Design of Scheduling Policy Extender
The Scheduling Policy Extender provides a customizable policy interface and a utility function module designed to facilitate the rapid prototyping and evaluation of scheduling policies. By extending the provided interfaces and leveraging utility functions, users can efficiently develop tailored scheduling strategies that meet the diverse performance, scalability, and reliability requirements of microservice-based applications in cloud-edge environments.
Figure 8. Overview of the Scheduling Policy Extender, highlighting extensible classes, utility functions, and example policies.
7.1. Policy Customization Interface
@@
The AbstractSchedulingPolicy also provides customizable internal methods such as single-pod scheduling (for single-service optimization), batch-pod scheduling (for multi-service optimization), and an update-metrics method triggered during re-scheduling events. By overriding these abstract methods, researchers and developers can incorporate sophisticated decision-making algorithms—ranging from heuristics to advanced machine learning techniques—while seamlessly integrating their logic with the underlying scheduling framework.
7.2. Utility Function Module
@@
7.3. Examples of Customized Policies
7.3.1. Policy 1: Call-graph–Aware
Microservice applications often exhibit dynamic call-graphs triggered by different request types, resulting in varying communication intensities between service pairs. This policy strategically co-locates microservices with high mutual communication on the same node or nodes with superior interconnectivity, thus minimizing cross-node traffic. To avoid potential node overload (e.g., excessive microservice instances in one node), stricter node capacity checks can be incorporated into the policy.
7.3.2. Policy 2: Latency-Aware
7.3.3. Policy 3: Bandwidth-Aware
Certain microservices generate substantial data transfers, making them bandwidth-sensitive. In scenarios with dynamic or constrained bandwidth conditions, this policy assigns high-bandwidth-demanding microservices to nodes offering the best available bandwidth or co-locates them to reduce inter-node data movement, thus optimizing throughput and performance.
7.3.4. Policy 4: Hybrid-dynamics-Aware
Considering the combined dynamics of varying call-graphs and fluctuating cross-node delays, this policy formulates the microservice scheduling as a Service-Node Mapping Problem. It strategically maps microservices to nodes by jointly optimizing the communication cost, which depends on both inter-service traffic volume and inter-node communication delays. Specifically, microservices with high interaction volumes are placed on nodes with low mutual communication delays. This integrated approach minimizes overall communication latency and maximizes performance under realistic, dynamic cloud-edge conditions.
8. Performance Evaluation
Table 1. Injected versus measured cross-node communication delays (ms). Each cell shows the injected delay (left) and the corresponding measured actual delay (right). The results demonstrate successful injection of varied delays from source nodes to multiple destinations, along with reserved channels without injected delays (i.e., Master node and google.com).
@@ 8. Performance Evaluation
Table 2. Saturated versus measured cross-node bandwidths (Mbits/sec). Each cell shows the saturated bandwidth first, followed by the actual measured bandwidth, which is emulated and measured by our proposed components in iDynamics.
@@
8.1. Cloud-edge Testbed Setup

### v3

@@ V-A1 Istio Service Mesh Implementation
To achieve fine-grained observation into bidirectional traffic between UM and DM microservices, including their replicas, we implemented the UM-DM Traffic Profiler using the Istio service mesh [^1] Istio is an open-source service mesh employing sidecar containers to facilitate secure communication, traffic management, and load balancing. [33]. The service mesh is an additional infrastructure layer integrated into containerized microservice deployments. With Istio deployed, all microservice traffic is transparently proxied via injected sidecar containers.
@@
In Kubernetes clusters with Istio service mesh enabled, each microservice pod includes the application container alongside an Envoy sidecar container. As illustrated in Figure 5, the service mesh architecture proxies traffic among microservice containers. In the Data Plane, orange rectangles represent the microservice containers, while green rectangles denote the sidecar containers that manage traffic and maintain dependencies between services. The green lines indicate data flows and dependency relationships among microservices. In the Control Plane, the Istio service mesh manages services and interactions via components including Pilot, Mixer, and Citadel for different purposes.
@@
Figure 5: Overview of a service mesh structure consisting of data plane and control plane.
Figure 6: A running example of a call graph with dependencies and traffic flows.
V-A2 Overhead Analysis
@@
Deploying a service mesh introduces additional processing delays due to the sidecar proxy layer. However, empirical evaluations indicate these delays minimally impact overall latency performance. According to [34], in Istio 1.21.2 with telemetry v2 enabled, each request passing through client-side and server-side Envoy proxies incurs a latency increase of approximately 0.18 milliseconds at the 90th percentile and 0.25 milliseconds at the 99th percentile compared to baseline performance without proxies. These results, obtained using a 1kB payload at 1000 requests per second across varying client connections (2, 4, 8, 16, 32, 64), were measured at the CNCF Community Infrastructure Lab [35]. Thus, the added latency overhead from the service mesh is negligible and acceptable for most practical scenarios.
@@
2:Output: Directed call-graph $G$ with weighted traffic edges
@@
V-B Call-graph Builder
We introduce the following terminology:
$\text{stress}^{\mu}_{\sigma}(\mu^{UM},\sigma^{DM},\Delta t)=\frac{\text{Bi-directional\_traffic}(\mu^{UM},\sigma^{DM})}{2\Delta t}$ (1)
[…]
Algorithm 1 constructs the call graphs with dependency and associated traffic load based on the defined stress elements. Initially, the algorithm retrieves all running microservices deployed within a namespace, which separates different applications. It then initializes a directed graph, representing each microservice as a graph node. The algorithm calculates the stress for each microservice pair using real-time traffic metrics collected over the specified interval. If the computed traffic stress exceeds zero, an edge representing the dependency and corresponding traffic load is created. The constructed graph thus reflects both the dependency topology and associated traffic stress, enabling accurate evaluation of scheduling decisions. Based on Algorithm 1, a running example for a representative microservice application (Social Network from [32]) is illustrated in Figure 6.
@@ V-B Call-graph Builder
Figure 7: (a) Packets are enqueued and dequeued according to classful disciplines. (b) Destination-specific delay emulation. (c) Destination-specific bandwidth shaping.
VI Design of Networking Dynamics Manager
The Networking Dynamics Manager in the proposed framework iDynamics consists of two main components: the Emulator and Measurer. In cloud-edge computing, accurately emulating dynamic network conditions such as latency variations and bandwidth fluctuations is critical for evaluating scheduling policies against SLA targets. However, performing these evaluations directly in production environments is challenging and inefficient. iDynamics addresses this issue by providing a controllable networking dynamics emulator that accurately replicates delays and bandwidths. Moreover, precisely and efficiently measuring these emulated dynamics is essential for reliable evaluation.
VI-A Linux Traffic Control
@@
u32 filter: Filters in tc play a crucial role in directing traffic to specific classes or queues by identifying and matching packet characteristics. u32 filter is one of the most versatile and widely used filters in tc. It allows users to match arbitrary bits of a packet header, enabling particular rules. For example, users can match fields like source/destination IPs, port numbers, and protocol types by extracting specific bits from the header.
For the dynamic network emulations in iDynamics, classful qdiscs and u32 filter are implemented to meet the goals of changing cross-node communication delays and bandwidths in the cloud-edge continuum. Moreover, after injecting the dynamic delays or shaping available bandwidths, Cross-node Measurer is designed to measure these delays and bandwidths in the current cloud-edge continuum because the accurate and efficient measurement of these dynamics is crucial for developing and evaluating networking-related scheduling policies.
@@
VI-B Emulator: A classful qdisc design
To address these limitations, we propose a customized delay injection scheme that classifies packets using filters to distinguish egress packets and directs them based on their IP destinations (see Figure 7b). Additionally, we reserve an extra channel for default packet transmission without injected delays, ensuring that the performance of other services remains unaffected, for example, the packets from/to the master node and outside internet services like Google. We implemented this scheme using the tc networking tool and the Hierarchical Token Bucket (htb).
Furthermore, we employ the technique illustrated in Figure 7 to design an algorithm for emulating customized cross-node delays, as presented in Algorithm 2. To emulate cross-node delays more realistically, we incorporate several factors: the base latency bl (representing the ideal minimal latency); the maximum additional latency mal; the distance factor df (accounting for latency due to physical separation); and the congestion factor cf (which simulates network uncertainties induced by traffic congestion). Here, $\textsc{RandUni}(0,mal)$ denotes the additional delay generated according to a random uniform distribution between 00 and $mal$. Thus, the emulated communication delay from node $i$ to node $j$ in a cluster of $N$ cloud-edge nodes is given by the following equation:
$\text{delay}^{j}_{i}=\left\lfloor\left(bl+\textsc{RandUni}(0,mal)\times\frac{\lvert i-j\rvert}{N}\right)\times cf\right\rfloor$ (2)
@@ VI-B Emulator: A classful qdisc design
Figure 8: Design of Measurer for delay and bandwidth measurement through distributed agents across cloud-edge nodes.
Emulation for Customized Bandwidths: Shaping customized available bandwidths between different cloud-edge node pairs is also crucial for creating dynamic networking conditions, thereby enhancing the generalizability of evaluated scheduling policies. Prior research—such as the work presented in [21] and [25]—has demonstrated the feasibility of emulating dynamic network conditions, including bandwidth variations, using decentralized emulation techniques. However, these approaches typically apply a global or node-level configuration that does not allow fine-grained, per-destination control.
In contrast, our proposed emulation scheme leverages the classful qdisc with unique traffic classes and utilizes the u32 filter to match different IP destinations. As illustrated in Figure 7c, this approach enables the emulation of distinct bandwidth settings for different destination nodes, even when originating from the same source cluster node. This fine-grained control is essential for accurately emulating heterogeneous network conditions in cloud-edge environments.
VI-C Measurer: Distributed Agent Design
In practical cloud-edge computing environments, fluctuating network conditions significantly impact microservice performance and SLA compliance, including varying delays and bandwidths between nodes. High latency between cluster nodes negatively affects microservice communication, while dynamically changing bandwidth can lead to congestion, increased packet loss, and potential SLA violations. To accurately capture these dynamic cross-node conditions, we introduce the Measurer, a distributed agent-based measurement module within the Networking Dynamics Manager component of iDynamics (Figure 4).
@@ VI-C Measurer: Distributed Agent Design
Design Overview: The Measurer uses a unified approach for capturing cross-node delays and bandwidths through distributed measuring agents. Although delay and bandwidth measurements follow a similar structure, we illustrate the design using delay measurement for clarity. The module consists of a centralized information processing unit and distributed lightweight agents. The centralized processing unit maintains minimal connections with agents, aggregating the collected data efficiently.
@@
The distributed agents run as lightweight containers deployed across cluster nodes, dedicated to measuring and reporting network metrics. We implement an automatic scaling mechanism for these agents to ensure robustness and adaptability during cluster scaling. When new nodes join the cluster, corresponding agents are automatically instantiated, ensuring consistent and accurate measurements. Conversely, when nodes are removed, their associated agents are gracefully terminated.
@@
Figure 9: Overview of the Scheduling Policy Extender, highlighting extensible classes, utility functions, and example policies.
Overhead Analysis: Regarding the overhead of measurement in a real cloud-edge cluster, we optimized the measurement using parallel processing. The measurement tasks for delays and bandwidths are designed to run as concurrent tasks, and each of the measured results is aggregated into a result dictionary. Additionally, the distributed delay measuring agents are lightweight (about 0.2 MiB) and stable (running over 6 months without any failure). Each node runs a delay-measuring agent developed from the curlimages/curl image, consuming around 0.2 MiB of memory per node. Similarly, each node also runs a bandwidth-measuring agent implemented from the networkstatic/iperf3 image, consuming around 0.84 MiB of memory on each node. As the measurement tasks are designed to run in parallel as well, the measuring speed can be adaptively tuned by increasing or decreasing the concurrency, depending on the current load levels in the cluster.
VII Design of Scheduling Policy Extender
The Scheduling Policy Extender provides a customizable policy interface and a utility function module to facilitate the rapid prototyping and evaluation of scheduling policies. By extending the provided interfaces and leveraging utility functions, users can efficiently develop tailored scheduling strategies that meet microservice-based applications’ diverse performance, scalability, and reliability requirements in cloud-edge environments.
VII-A Policy Customization Interface
@@
The AbstractSchedulingPolicy also provides customizable internal methods such as single-pod scheduling (for single-service optimization), batch-pod scheduling (for multi-service optimization), and an update-metrics method triggered during re-scheduling events. By overriding these abstract methods, researchers and cloud practitioners can incorporate sophisticated decision-making algorithms—ranging from heuristics to advanced machine learning techniques—while seamlessly integrating their logic with the underlying scheduling framework.
VII-B Utility Function Module
@@
TABLE II: Injected versus measured cross-node communication delays (ms). Each cell shows the injected delay (left) and the corresponding measured actual delay (right). The results demonstrate successful injection of varied delays from source nodes to multiple destinations, along with reserved channels without injected delays (i.e., Master node and google.com).
@@
TABLE III: Saturated versus measured cross-node bandwidths (Mbits/sec). Each cell shows the saturated bandwidth first, followed by the actual measured bandwidth, which is emulated and measured by our proposed components in iDynamics.
@@
VII-C Examples of Customized Policies
VII-C1 Policy 1: Call-graph–Aware
Microservice applications often exhibit dynamic call graphs triggered by different request types, resulting in varying communication intensities between service pairs. This policy strategically co-locates microservices with high mutual communication on the same node or nodes with superior interconnectivity, thus minimizing cross-node traffic. To avoid potential node overload (e.g., excessive microservice instances in one node), stricter node capacity checks can be incorporated into the policy.
VII-C2 Policy 2: Latency-Aware
VII-C3 Policy 3: Bandwidth-Aware
Certain microservices generate substantial data transfers, making them bandwidth-sensitive. In dynamic or constrained bandwidth conditions, this policy assigns high-bandwidth-demanding microservices to nodes offering the best available bandwidth or co-locates them to reduce inter-node data movement, thus optimizing throughput and performance.
VII-C4 Policy 4: Hybrid-dynamics-Aware
Considering the combined dynamics of varying call graphs and fluctuating cross-node delays, this policy formulates the microservice scheduling as a Service-Node Mapping Problem. It strategically maps microservices to nodes by jointly optimizing the communication cost, which depends on both inter-service traffic volume and inter-node communication delays. Specifically, microservices with high interaction volumes are placed on nodes with low mutual communication delays. This integrated approach minimizes communication latency and maximizes performance under realistic, dynamic cloud-edge conditions.
TABLE IV: Performance comparison between Kubernetes default scheduling and Policy 1 (Call-graph–Aware) under varying QPS and call-graph dynamics. Metrics include average (Avg), 99th percentile (p99), and standard deviation (Stdev) of response time (ms). SLA violations are highlighted in red, and SLA-compliant improvements in green.
Policy Evaluation (5 Cluster Nodes) Running time 0–8 min (Call-graph 1) 8–16 min (Call-graph 2) 16–24 min (Call-graph 3) 24–32 min (Call-graph 4, mix)
QPS (req/s) 30 10 50 70 30 10 50 70 30 10 50 70 30 10 50 70
Kubernetes Policy (default) Avg 34.96 36.07 34.97 34.64 188.5 191.8 191.8 194.4 80.07 77.36 73.68 68.85 63.08 63.07 60.58 59.84
p99 63.49 70.33 62.40 61.73 336.7 269.3 349.7 373.3 202.1 202.2 200.9 198.9 232.2 281.86 222.3 218.4
Stdev 6.020 5.930 12.75 10.27 27.22 18.80 36.97 44.79 45.80 40.86 37.75 30.88 53.05 56.06 49.14 48.12
Call-graph-Aware (Policy 1) Avg 34.74 35.84 34.26 34.30 188.6 63.59 41.54 40.96 46.03 45.87 42.74 41.12 30.00 32.47 29.43 29.35
p99 60.22 67.90 60.22 60.00 338.4 463.6 73.92 73.47 182.4 182.0 176.0 159.9 104.5 114.7 109.9 104.5
Stdev 4.220 4.790 4.120 4.560 27.04 75.87 6.100 6.160 40.88 38.50 35.07 30.05 18.38 20.85 18.54 18.48
(a) 5 Cluster Nodes (SLA violations (>100 ms))
Policy Evaluation (10 Cluster Nodes) Running time 0–8 min (Call-graph 1) 8–16 min (Call-graph 2) 16–24 min (Call-graph 3) 24–32 min (Call-graph 4, mixed)
QPS (req/s) 30 10 50 70 30 10 50 70 30 10 50 70 30 10 50 70
Kubernetes Policy (default) Avg 13.27 14.36 13.22 13.22 168.6 171.8 171.2 174.2 57.33 56.06 49.92 46.11 40.67 44.15 39.31 40.15
p99 20.24 38.91 18.08 18.64 299.1 243.8 315.6 326.9 197.8 203.1 197.3 187.5 207.7 259.7 203.5 203.5
Stdev 2.460 4.610 2.510 1.880 24.87 20.17 33.17 39.46 48.30 45.73 38.64 31.42 51.43 60.34 49.63 50.49
Call-graph-Aware (Policy 1) Avg 13.89 14.84 13.51 13.33 168.88 127.6 117.26 118.4 71.32 70.87 68.40 61.83 34.92 39.37 35.20 35.69
p99 23.85 42.72 17.38 20.00 300.8 594.4 210.7 218.75 309.8 314.4 310.0 305.4 202.2 251.3 183.4 194.3
Stdev 3.260 5.260 2.100 2.130 25.27 92.80 19.12 22.33 68.42 68.38 63.62 49.84 43.76 51.00 43.75 43.88
(b) 10 Cluster Nodes (SLA violations (>150 ms))
Policy Evaluation (15 Cluster Nodes) Running time 0–8 min (Call-graph 1) 8–16 min (Call-graph 2) 16–24 min (Call-graph 3) 24–32 min (Call-graph 4, mix)
QPS (req/s) 30 10 50 70 30 10 50 70 30 10 50 70 30 10 50 70
Kubernetes Policy (default) Avg 41.31 43.12 40.83 40.50 231.4 234.45 236.1 241.4 64.34 65.07 62.22 60.61 71.19 74.00 68.90 67.99
p99 92.10 103.5 89.21 78.14 425.7 360.9 445.4 492.5 98.88 96.64 104.9 108.0 302.8 384.7 278.0 277.8
Stdev 7.36 10.40 7.27 6.68 38.79 29.23 53.02 62.90 12.41 11.55 11.92 11.69 116.0 74.33 61.18 60.84
Call-graph-Aware (Policy 1) Avg 41.32 43.25 40.83 40.56 231.6 151.4 129.8 130.4 76.48 70.24 67.46 64.76 57.69 60.89 55.56 55.78
p99 91.84 98.24 88.77 78.91 418.6 364.5 243.2 240.4 345.6 349.7 325.9 309.8 239.6 293.9 222.1 223.7
Stdev 7.670 10.19 6.990 7.310 37.46 101.7 25.57 26.73 79.79 71.62 65.01 58.29 44.53 51.83 40.55 40.90
(c) 15 Cluster Nodes (SLA violations (>200 ms))
VIII Performance Evaluation
VIII-A Cloud-edge Testbed Setup